### v1

# Towards the Pedagogical Steering of Large Language Models for Tutoring: A Case Study with Modeling Productive Failure

Abstract
One-to-one tutoring is one of the most efficient methods of teaching. Following the rise in popularity of Large Language Models (LLMs), there have been efforts to use them to create conversational tutoring systems, which can make the benefits of one-to-one tutoring accessible to everyone. However, current LLMs are primarily trained to be helpful assistants and thus lack crucial pedagogical skills. For example, they often quickly reveal the solution to the student and fail to plan for a richer multi-turn pedagogical interaction. To use LLMs in pedagogical scenarios, they need to be steered towards using effective teaching strategies: a problem we introduce as Pedagogical Steering and believe to be crucial for the efficient use of LLMs as tutors. We address this problem by formalizing a concept of tutoring strategy, and introducing StratL, an algorithm to model a strategy and use prompting to steer the LLM to follow this strategy. As a case study, we create a prototype tutor for high school math following Productive Failure (PF), an advanced and effective learning design. To validate our approach in a real-world setting, we run a field study with 17 high school students in Singapore. We quantitatively show that StratL succeeds in steering the LLM to follow a Productive Failure tutoring strategy. We also thoroughly investigate the existence of spillover effects on desirable properties of the LLM, like its ability to generate human-like answers. Based on these results, we highlight the challenges in Pedagogical Steering and suggest opportunities for further improvements. We further encourage follow-up research by releasing a dataset of Productive Failure problems and the code of our prototype and algorithm.
Figure 1: Schematic representation of StratL (in blue), an algorithm to control an LLM’s tutoring strategy.
1 Introduction
One-to-one tutoring is an effective teaching method. Because of the efficiency of tutoring and its prohibitive cost, the challenge of finding affordable one-to-one tutoring methods is considered as one of the greatest problems in education: the so-called Two-Sigma Problem (Bloom 1984; Cohen, Kulik, and Kulik 1982). Following the rising popularity of Large Language Models (LLMs), there has been a growing interest in using them to make tutoring affordable by building conversational tutoring systems (Jurenka et al. 2024; Khan 2023) However, research empirically pointed out the lack of pedagogical properties of LLMs (Macina et al. 2023b; Tack and Piech 2022). These problems stem from the fact that most LLMs are instruction-tuned by optimizing for human preferences (Ouyang et al. 2022; Zhang et al. 2023). As a result, LLMs typically provide assistant-like responses. Yet a tutor’s goal is to maximize learning, not user satisfaction.
These two goals can be contradictory. For example, if a student asks for help on a problem, directly providing them with the solution would likely maximize their satisfaction, and be the answer preferred by LLMs (Macina et al. 2023a). Yet, it is more efficient from a learning standpoint to promote students’ active engagement to let them cognitively engage with the problem, for example by providing an indirect hint or by asking a question (Freeman et al. 2014; Merrill et al. 1992). Moreover, existing LLM tutors are optimized and mostly evaluated for single-turn requests, while a tutoring dialog should be envisioned as a multi-turn teacher-student interaction where the teacher uses various pedagogical cues (Merrill et al. 1992). As LLMs are not natively made for tutoring, we introduce the fundamental shift from their user-serving goal to a pedagogically suitable one as the Pedagogical Steering problem.
We present and evaluate StratL, an algorithm to address the Pedagogical Steering problem. To correct for a pedagogically suitable goal, we formally model the concept of a multi-turn tutoring strategy. We define a multi-turn strategy as a succession of single-turn pedagogical goals called tutoring intents. StratL dynamically redefines the tutoring intents after every student utterance using a transition graph and prompts the LLM to follow them for the next turn. We partner with learning scientists to integrate a strategy based on an advanced learning design known as Productive Failure (PF) (Kapur 2008). A meta-analysis of 53 studies showed that PF is more effective for human learning than other classical teaching methods (Sinha and Kapur 2021). PF can be seen as the most extreme version of not directly telling the answer to the students, as it intentionally lets them explore the problem and confront them with various suboptimal solutions, for example by using Socratic questioning. While PF works across learning subjects, this work focuses on high school math, as it has been most explored in PF research (Sinha and Kapur 2021), and prototypes ablations of StratL to evaluate it in real-world settings.
We evaluate StratL and uncover practical limitations and insights on the Pedagogical Steering problem through two studies: a preliminary study with simulated students and a field study with 17 high school students in Singapore. Together, these studies show that StratL effectively steers the LLM to adhere to the Productive Failure tutoring strategy and does so with no negative spillover effects on other desirable properties of the LLM.
We summarize our contributions as follows:
  • We define the problem of Pedagogical Steering in the context of LLM multi-turn tutoring. We explain why instruction-tuned LLMs are misaligned for tutoring from a Learning Sciences standpoint, and review the current state of research from this perspective.
  • We introduce the StratL algorithm to model the concept of tutoring strategy and to prompt an LLM towards a given strategy using single-turn intents.
  • We run a field test in a Singaporean high school to evaluate the potential real-world impact and gather insights on the Pedagogical Steering problem. We measure the extent to which StratL and its ablations control the pedagogy followed by the LLM through Learning Sciences–inspired metrics and investigate the limitations.
  • We release a dataset of Productive Failure tutoring exercises. These exercises are particularly interesting because each has multiple suboptimal solutions and students often require multiple interactions to find the optimal one.
2 Background and Related Work
Pedagogical shortcomings of LLMs: Pedagogical Steering problem
Recent advances enable state-of-the-art LLM-based systems to reliably perform middle– to high school–level mathematics, such as using prompting (Wu, Zhang, and Huang 2023; Kojima et al. 2022), multiple sampling strategies with consistency sampling (Wang et al. 2023) or refinement (Madaan et al. 2024; Shridhar et al. 2023; Li et al. 2023), fine-tuning (Zelikman et al. 2022; Yuan et al. 2023) or use of external tools like symbolic solvers or code interpreters (He-Yueya et al. 2023; Mialon et al. 2023). These reasoning capabilities could give LLMs the ability to serve as a basis for lesson-agnostic tutors.
Notwithstanding the promising capabilities of LLMs, several studies exhibited their poor pedagogical properties (Tack and Piech 2022). Among other flaws, most models often reveal the solution of the exercise to the students, even if prompted not to do so (Macina et al. 2023a). Their inefficiency as learning tools has also been measured more directly: when comparing students who were asked to solve problems using GPT-generated hints or human-generated ones, the human-generated hints group achieved substantially and statistically significantly higher learning gains (Pardos and Bhandari 2023).
These pedagogical limitations stem from a core problem: LLMs are not optimized for tutoring. At a fundamental level, autoregressive LLMs are transformer-based models trained to take a sequence of tokens as input and predict the most likely following one (Zhao et al. 2023). After this initial pre-training process, an alignment step is necessary to turn these base models into useful assistants able to follow human instructions, like ChatGPT. The alignment is often performed by fine-tuning LLMs on human preferences (Zhang et al. 2023), mainly using a framework known as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al. 2022). During this process, LLMs learn to comply with requests by optimizing user satisfaction.
LLM assistants thus generate texts that optimize for human preferences. This is misaligned with the way a tutor should behave. Indeed, when a student asks for help, the goal of the tutor should not necessarily be to maximize the student’s satisfaction, but rather their learning. For this, the tutor may have an intent (for example, asking back a question) different from what the student would like them to do (for example, being given the solution). This mismatch is one of the core limitations of LLMs as tutors.
LLMs for Tutoring
Several works have attempted to address the lack of pedagogical properties of LLMs. The easiest and most direct approach to address this problem is to write an extensive prompt (Wang et al. 2024; Sonkar et al. 2023; Chowdhury, Zouhar, and Sachan 2024). However, enumerating all desired behaviors of a tutor in words is complex, especially for more nuanced pedagogical strategies.
Another method considered is fine-tuning LLMs to efficient tutoring strategies. However, supervised fine-tuning is hampered by the scarcity of tutoring datasets, as their collection is costly and poses ethical concerns (Kasneci et al. 2023). Macina et al. (2023a) adopt an inverted Wizard-of-Oz approach by asking human teachers to tutor GPT-simulated students to create a comprehensive dataset of high-quality tutoring sessions and show that fine-tuning LLMs on this dataset improves their pedagogy. Similarly, the CLASS framework (Sonkar et al. 2023) uses an LLM to generate a synthetic conversational dataset with student-tutor interactions to fine-tune a smaller LLM. However, the quality of these conversations is not evaluated, and is prone to hallucinations. Moreover, fine-tuning is expensive and restricts the LLM to the specific teaching style present in the training data, thus requiring multiple datasets for different pedagogical approaches. The Bridge method (Wang et al. 2024), which we use as an inspiration for one of our baselines, uses an LLM to classify student errors and decide tutor interventions on the conversation history. Systems based on Bridge or CLASS let the LLM choose the tutoring strategy on its own, assuming its ability to make pedagogically sound choices. In contrast, our approach focused on letting learning scientists precisely control it. To do so, we employ a structured pedagogical intent selection using a transition graph based on classified student states, enabling experts to constrain the model to more faithfully follow a given pedagogy over multi-turn conversations.
Recent work reconsiders the CLASS framework from the point of view of alignment and tackles it using common alignment techniques such as RLHF (Sonkar et al. 2024). However, the data are synthetically generated which leads to questionable quality and bias, and as for fine-tuning, the need for data is the main drawback of this technique.
Productive Failure (PF)
In education, “Instruction followed by Problem Solving” approaches, which directly instruct students on how to approach a problem and then let them practice, are widely used. However, Productive Failure (Kapur 2008, 2016) is a “Problem Solving followed by Instruction” learning design in which the learner attempts to solve open-ended problems before formally learning how to approach them, to build a deeper understanding of the topic. With PF, the teacher uses mostly questioning to let students cognitively engage with the problem. There is a large body of evidence that PF is effective with meta-analyses reporting an effect size of 0.87 standard deviations (Sinha and Kapur 2021), high by the learning sciences standards, and performing better than directly instructing students or telling them solutions.
A successful problem-solving session in PF provides opportunities for the learner to explore the problem and build as many Representation and Solution Methods (RSMs) as possible, even if they fail at solving the problem (Kapur and Bielaczyc 2012). In simple terms, these RSMs can be thought of as mental representations of the problem, or partial solutions. The number of RSMs generated is one of the best predictors of how much students learn from PF (Kapur and Rummel 2012). Our tutoring strategy thus aims to promote the generation of such RSMs, and we will use their number as one of the evaluation metrics.
Overall, PF is a suitable learning design to validate our approach because (1) it is a widely successful pedagogical strategy, and (2) opposed to native LLM behavior as a PF tutor should not directly reveal the solution to the students.
3 Methodology
3.1 Formalism
Dialog Tutoring Task
We follow the formalization of the Dialog Tutoring task defined in (Macina et al. 2023b). We consider a tutoring dialog history $\mathcal{H}_{\leq n}=(u_{1},\ldots,u_{n})$ made of tutoring turns $u_{t}=(y_{t},x_{t})$ with $y_{t}$ the tutor’s utterance and $x_{t}$ the student’s utterance at turn $t$. $x_{t}\in\mathcal{V}^{\ast}$ and $y_{t}\in\mathcal{V}^{\ast}$ are finite sequences of tokens from a vocabulary $\mathcal{V}$. Each tutor utterance is associated with a set of tutoring intents $\mathbf{i}_{t}$ whose elements are chosen from a taxonomy of intents =$\mathcal{I}=\{I_{1},\ldots,I_{|\mathcal{I}|}\}$. An intent is defined as a goal of the tutor utterance, e.g. ‘correct the student’s mistake’. We denote by $K$ the grounding information, i.e. additional information provided to the LLM to give context. In our case, $K$ includes the tutoring exercise and its solution. The resulting generative conditional model is:
$y_{n+1}\sim p(\ y\ |\ \mathbf{i}_{n+1},\mathcal{H}_{\leq n},K)$ (1)
Tutoring Strategy Modeling
We define a tutoring strategy predictor as a function $\phi$ that takes as input $(\mathcal{H}_{\leq t},K)$, the grounding and dialog history up to any turn $t$, and that outputs a set $\mathbf{i}_{t+1}\subseteq\mathcal{I}$ of single-turn tutoring intents for turn $t+1$. Abstractly, $\phi$ represents the decision process by which an expert tutor, given the context of the conversation, selects the best possible intents from a pedagogical standpoint to react to the student’s last utterance. The resulting conditional tutor model is:
$y_{n+1}\sim p(\ y\ |\ \phi(\mathcal{H}_{\leq n},K),\mathcal{H}_{\leq n},K)$ (2)
3.2 Tutoring Strategy – Learning Sciences Intuitions
Expert tutors base their interventions on observations about (1) the student’s emotional or knowledge state (Lepper et al. 1993), and (2) the state of the conversation and their previous interventions (Merrill et al. 1992). For example, a tutor should not react the same way to a computational mistake than to a reasoning one. Based on these intuitions, we suggest further refining the high-level intent selection function $\phi$, which abstractly depends on the dialog history and grounding, to a more practical function $\widetilde{\phi}$ that selects the next intents based on state features (1) and the previous intents (2). We propose the formalization developed in the next paragraph.
Student State Tracing
Each turn is associated with a set $\mathbf{s}_{t}$ of state features describing the state of the student and of the conversation at turn $t$. The state features are chosen from a set $\mathcal{S}$ of possible features. A feature in $\mathcal{S}$ could, for example, be ‘the student is confused’ or ‘the student made a computational mistake’. We denote $\psi$ the abstract function that provides these features based on the dialog history and grounding. We approximate $\phi$ with a function $\widetilde{\phi}(\mathbf{i}_{n},\mathbf{s}_{n})=\mathbf{i}_{n+1}\subseteq\mathcal{I}$. $\widetilde{\phi}$ depends on the student state features $\mathbf{s}_{n}$ and previous tutor intents $\mathbf{i}_{n}$. Our final conditional generation model is:
$y_{n+1}\sim p(\ y\ |\ \widetilde{\phi}\left(\mathbf{i}_{n},\psi(\mathcal{H}_{\leq n},K)\right),\mathcal{H}_{\leq n},K)$ (3)
3.3 Algorithm
The Pedagogical Steering problem is caused by a misalignment between an LLM’s goal and tutoring goals, and by LLMs’ single- rather than multi-turn dimension. To steer the LLM towards a multi-turn pedagogical strategy, we introduce StratL, an algorithm that dynamically updates the tutoring intents after every student utterance and prompts the LLM to follow them at the single-turn level.
A visual representation of StratL is presented in Figure 1. StratL is an algorithm for the formalization given in Section 3.1. This algorithm is the succession of three procedures:
  1. I.
    Student State Tracing
  2. II.
    Tutor Intent Selection
  3. III.
    Intent Dependent Steering
I. The Student State Tracing mirrors the function $\psi$. This step is a multi-label classification of the dialog history, and the latest student utterance at turn $t$. The selected labels (called features) $\mathbf{s}_{n}$ are chosen from a pre-defined set of possible features $\mathcal{S}$. We defined features that provide the most relevant information for the subsequent Intent Selection. These features broadly fall into the categories of (1) error types, (2) student requests, (3) states of conversation, and (4) emotional states. These feature categories each provide valuable information to be used in the Intent Selection. (1) helps plan error remediation (e.g. a reasoning and a computational mistake should not be dealt with equally) (Daheim et al. 2024; Wang et al. 2024), (2) is necessary as different student requests call for different reactions (e.g. if the student asks for the solution v.s. if they ask for a definition), (3) helps structure the conversation (e.g. knowing if the problem is solved) and (4) allow for proper emotional responses from the tutor (Lepper et al. 1993). The precise features we use are detailed in Appendix A.2. State tracing is also performed by an LLM prompted to identify these different features given the set of possible student states $\mathcal{S}$ and the dialog history. Implementation details such as the prompt we used are provided in Appendix A.2.
II. The tutor intent selection mirrors the function $\widetilde{\phi}$. We model tutor intent selection as a deterministic transition from the previous set of intents $\mathbf{i}_{n}$ to the next $\mathbf{i}_{n+1}$, conditioned on the output $\mathbf{s}_{n}$ of step I. This step can be represented by a transition graph that shows how to move from one intent to the other depending on the selected features. This graph defines the tutoring strategy $\widetilde{\phi}$. We collaborated with Learning Scientists to manually design a transition graph mirroring a PF tutoring strategy. The graph can be found in Appendix A.3. The transition rules can be defined by an expert, as we did, but we also encourage future work to learn them from annotated data. We used the theories of Scaffolding (Reiser 2004), Affective Cognition (Lepper et al. 1993) and Productive Failure (Sinha and Kapur 2019; Kapur and Bielaczyc 2012) to design a taxonomy of tutoring intents $\mathcal{I}$. This taxonomy is detailed in Appendix A.1. These intents are mainly separated into (1) Scaffolding intents (such as giving a hint) that help the student by restricting the solution space, and (2) Problematizing intents (such as asking to formalize an intuition) that help the student deepen their understanding by exploring the solution space.
For our Productive Failure strategy, we prioritize light Scaffolding intents when the student is stuck and use Problematizing intents whenever the student shows an understanding of the problem, to encourage the free exploration of the problem and the generation of multiple RSMs. The exact Intent Selection transition graph we use is displayed in Figure 4 of Appendix A.3. A detailed explanation of the Learning Science principles we used to design our Productive Failure Intent Selection is given is Appendix A.5.
III. The Intent Dependent Steering generates the prompt $p_{n+1}$ for the LLM Tutor conditioned on the selected tutoring intents $\mathbf{i}_{n+1}$. This prompt integrates each intent in $\mathbf{i}_{n+1}$ by concatenating intent-dependent prompt additions to the LLM’s base prompt. We define a one-to-one correspondence between each possible tutoring intent in $\mathcal{I}$ and a prompt addition. This correspondence is given in Appendix A.4. As depicted in Figure 1, the generated prompt $p_{n+1}$ is made of a base prompt template, followed by the grounding information, and the intent-specific prompt additions.
4 Experiments
Experimental Setup. As a case study, we use StratL to implement a prototype high school math tutor based on GPT-4o (gpt-4o-2024-08-06) and following a PF–inspired tutoring strategy. Our tutor discusses one problem per tutoring dialog and has access to its solution. The precise implementation of the three steps of StratL in our prototype, including (I) the detailed prompt of the Student State Tracing and the set of possible state features $\mathcal{S}$, (II) the transition rules for the Intent Selection and the set of possible intents $\mathcal{I}$, and (III) the correspondence between intents and prompt additions, are given in Appendix A.2. We also detail the model specifications and parameters.
4.1 Research Questions
We validate the capacity of our system to control the tutoring strategy employed by the LLM by assessing the capacity of our prototype to follow the PF paradigm. We want our system to do so without hindering the factual correctness or desirable conversational properties of LLMs such as their coherence. We formulate these goals in two research questions:
  1. RQ1. To what extent does StratL make the LLM Tutor follow Productive Failure?
  2. RQ2. Does StratL influence on other desirable properties of the LLM Tutor?
    1. RQ2.1. Does StratL influence the ability of the LLM to generate factually correct responses?
    2. RQ2.2. Does StratL influence the ability of the LLM to act like a human?
    3. RQ2.3. Does StratL influence the perceived usefulness of the tutor?
In our implementation, We also chose to handcraft the Intent Selection mechanism with the help of PF expert teachers, with the assumption that an LLM could not model such a complex strategy by choosing intents on its own. This is different from Wang et al. (2024) which lets an LLM choose how to react. We thus formulate a design question to test our assumption:
  1. DQ1. Can the intent selection be left to an LLM?
4.2 Dataset
We collected a dataset of 20 problems from Productive Failure research. These problems are unique because each has multiple possible suboptimal solutions (and a unique optimal one) and requires multiple interactions from students to explore their solution space. As such, they are high-quality testing problems to evaluate the pedagogical skills of a tutoring system.
We test our system on two problems from our dataset. Both are suitable for 9th-grade students. The problem Consistency is adapted from Kapur and Toh (2015) and is one of the most studied PF problems in the literature. This problem is an open-ended invention problem which asks students to create a measure of consistency (canonically measured by the variance) to determine the most consistent football player among three based on the number of goals they scored over the past years. An effective PF tutor for this problem would help the students construct as many ways to measure consistency as possible. The second problem, entitled Country, is an open-ended problem taken from Kerrigan (2018) which asks students to find a way to fairly distribute voting seats for the different states of a country based on their population. For this problem, a good PF tutor would help the student without giving away any of the crucial reasoning steps necessary to solve this problem, such as using proportionality. The exact problem statements, their corrections, and the dataset specification are provided in Appendices A.6 and A.7.
4.3 Baselines
We test our StratL-controlled tutor LLM against three ablation versions.
  • V1:
    Full StratL. An LLM controlled by our implementation of StratL.
  • V2:
    No Intent. Baseline version leaving complete freedom to the tutor LLM to choose the best way to answer without conditioning on any intent. We use it to assess the tutoring strategy fidelity of our system.
  • V3:
    Constant Intent. This version constantly uses the same intent (labeled “Seek Strategy” in Table 3 of Appendix A.1) that only asks the student to proceed with the exercise. We use this baseline to assess the importance of a well-crafted Intent Selection procedure.
  • V4:
    LLM Intent. In this version, the choice of the intent is left to an LLM similarly as in the related work (Wang et al. 2024). We use it to assess whether the intent’s choice needs to be determined by an expert-designed procedure or if a natural language description of the tutoring strategy to follow is enough (DQ1). Its prompt is given in Appendix A.10.
4.4 Evaluation Metrics
To quantitatively assess the Productive Failure fidelity of the different versions (their success in following PF), we have to rely on proxy variables. We propose as a main proxy the number of student-generated Representation and Solution Methods (RSMs, which can be thought of as partial solutions). As explained in Section 2, the diversity of student-generated RSMs is the first predictor of learning gains with PF (Kapur and Bielaczyc 2012; Kapur and Rummel 2012). We thus use this number to measure the PF fidelity of the tutor.
However, not every problem easily admits multiple RSMs. In our case, only Consistency does, as an invention problem. We thus define another proxy, the Productive Failure Score. One of the core ideas of PF is to encourage the student to construct their own understanding of the problem by exploring it by themselves, even if they do not succeed in solving it. Building on this idea, we define for each problem a grading rubric used to compute the tutor’s Productive Failure Score: For each problem, we identify a list of critical reasoning steps that represent different conceptual understandings of the problem. To assess a tutoring conversation, we give a point to the tutor for every crucial reasoning step that the student found on their own. We do not award the point if the tutor reveals it or directly hints to it. This metric can be thought of as an adjusted sub-step success rate. An efficient PF tutor would help the student explore the problem and thus achieve these conceptual understandings, without giving them any direct hint or revealing them, resulting in a high PF Score. We display the precise sub-steps for both testing problems in Appendix A.8.
4.5 Preliminary Study with Simulated Students
Because of ethical concerns (Kasneci et al. 2023), we first tested our prototype with GPT-4–simulated students before exposing it to humans. This method of using simulated students, known as interactive tutoring, is commonly employed in the literature (Sonkar et al. 2023; Macina et al. 2023a). Importantly, we do not assume that the LLM is able to accurately model students’ learning, especially given the somewhat counter-intuitive aspect of PF. However, this assumption is not necessary for our preliminary study as we are interested in the ability of our algorithm to follow a PF tutoring strategy, rather than the learning gains it produces. We simulated three conversations for each of the two problems and of the four versions of the tutor (V1–4). We use a scoring rubric to manually annotate the conversations to compute the PF Score and number of RSMs as described in Section 4.4.
4.6 Field Test with Students
We ran the field test with 17 9th-grade students from a high school in Singapore. As we explain in Section 5.1, we did not include V3 and V4 in the field test following the preliminary study results. Each student was randomly assigned to one of two conditions (V1 or V2), and they had 15 minutes to interact with the tutor. We manually annotated their conversations using our rubric to compute the PF score and the number of generated RSMs.
Post-study Questionnaire. After the tutoring session, we asked each student about their experience with the tutor. The aim was to detect any spillover effect of StratL on the other desirable properties of the tutor LLM. Specifically, students were asked to rate on a 3-point Likert scale the following three aspects of their AI tutor:
  1. 1.
    The tutor is factually coherent in its answers.
  2. 2.
    The tutor answers with empathy like a teacher.
  3. 3.
    The tutor is helpful in solving the problem.
5 Results and Discussions
5.1 Preliminary Study with Simulated Students
Table 1: Productive Failure fidelity metrics for different tutor versions and different problems in the preliminary study with simulated students. Best values in bold.
# RSMs PF.S
Consistency Country
V1. StratL 2.34 3.67 4.00
V2. No Int. 1 1.33 0.44
V3. Constant Int. 1.34 4.00 4.00
V4. LLM Int. 1.34 3.67 3.11
DQ1.
Using an LLM prompted with a natural language description of the tutoring strategy is not enough for the Intent Selection. The preliminary study in Table 1 suggests that StratL (V1) is successful in steering the LLM tutor towards PF, as evidenced by the greater number of RSMs elicited by this version. We observe that the version using StratL with an LLM responsible for the Intent Selection (V4) did perform better than the baseline LLM (V2), but not as well as StratL with our handcrafted Intent Selection mechanism. This result contributes to an answer to DQ1 by showing that providing a natural language description of the tutoring strategy to follow and letting an LLM choose the intents accordingly does not yield satisfactory results in terms of tutoring strategy fidelity.
RQ1.
The strategy followed by an LLM controlled by StratL is determined by the Intent Selection procedure. Overall, V3 achieves a rather high PF Score on the two problems, since it never provides any hint to students. However, this result is more a consequence of the way we defined the PF score rather than a sign of PF fidelity, as V3 does not perform as well on our second proxy: It does not manage to elicit as many RSMs from students as V1, with an average of only 1.34. This shows that the strategy followed by an LLM controlled by StratL is indeed determined by the Intent Selection procedure as it is the only part that differs between V1 and V3.
Based on these results, we decided not to include V3 and V4 in our field test with real students to focus on the best-performing version (V1) while keeping the non-StratL baseline (V2) for comparison.
Accuracy of Student State Tracing
The functioning of our system as a whole is contingent upon the functioning of the Student State Tracing. To test its reliability, we generate ten tutoring conversations for both problems and one author manually annotates them with ground truth state features $\mathbf{s}_{t}$. We compare the outputs of the Student State Tracing with the annotated data. As a multi-label classification task, we compute for each possible label (state feature) the precision, recall, and micro F1 score. We compare three different implementations of the State Tracing step and provide more details about our methodology and results in Table 6 of Appendix A.9. Our simple procedure already achieves satisfactory results of 0.77 for the micro F1 score. The Student State Tracing is a task of its own, and we encourage researchers to try different classification models to improve the reliability of this step.
Figure 2: Results from the field test and their confidence intervals. ‘PF.S’ refers to the ‘Productive Failure Score’ as defined in Section 4.4 and is out of 4. ‘RSMs’ denotes the number of student-generated RSMs. StratL (V1) succeeds in steering the LLM to follow Productive Failure as compared to V2.
5.2 Field Test with Students: Measuring end-to-end Effectiveness
RQ1. Tutoring Strategy Fidelity
StratL is successful in steering the LLM to follow the Productive Failure–inspired tutoring strategy. We report in Figure 2 the results of the protocol defined in Section 4.4. We note that the PF score (out of 4) is used as a proxy for the extent to which the tutor followed the PF tutoring strategy. We also report the average number of student-generated RSMs (partial solutions) on the problem Consistency. This metric is our best proxy for PF fidelity, as one of the goals of the PF tutoring strategy is to maximize this number, as explained in Section 2. Out of the two test problems, the number of RSMs can only be measured on Consistency as it is the only invention problem.
Table 2: Feedback questionnaire results (out of 3). ‘N’ refers to the number of students who answered the questionnaire, ‘Factual’ to the factual correctness of the tutor, ‘Human’ to its ability to communicate in natural language like a human, and ‘Helpful’ to its helpfulness as perceived by the student. We do not detect any harmful effect of StratL (V1) on these properties.
Consistency Country
N Factual Human Helpful N Factual Human Helpful
V1. StratL 5 1.80 1.60 1.20 3 3.00 2.67 2.67
V2. No Intent 4 2.25 1.75 2.00 5 3.00 2.20 2.80
[…]
RQ2. Student Feedback Questionnaire
We report in Table 2 the mean feedback questionnaire results for each problem on a 3-point Likert scale).
RQ2.1 StratL has no or little influence on the capacity of the tutor LLM to generate factually coherent responses. StratL (V1) scores similarly or slightly less than an LLM with no intent (V2) on factual correctness across the two problems. We do not detect any statistically significant difference for any of the versions.
RQ2.2 StratL does not influence the capacity of the LLM to generate human-like answers. None of the two versions is perceived as significantly more ‘human-like’ than the other. StratL does not seem to impede the ability of an LLM to produce natural language.
RQ2.3 StratL can influence the perceived helpfulness of the tutor LLM, through the tutoring strategy it makes it follow. We find that the baseline with no intents (V2) is perceived by the students as more helpful than V1. This is not surprising, as the PF tutoring strategy can be frustrating or misunderstood by students. Indeed, the PF strategy is designed to only give help when it is strictly necessary, which means that students spend more time being stuck and exploring suboptimal solutions. This can lead students to perceive the tutor as less helpful compared to a version that would give more help.
5.3 Qualitative Insights from Students Feedback
Our field test demonstrated the ability of StratL tutor LLM to follow Productive Failure. We highlight in this section some insights gathered from our students’ feedback.
Robustness
As discussed in Section 5.1, the Student State Tracing procedure can make labeling mistakes. These mistakes can then compromise the Intent Selection and result in the selection of contradictory intents. This can also happen if the Intent Selection is poorly designed. This is exemplified by the feedback of field study students: “The AI keeps on asking me to check my answer although it was correct”.
Grounding constraints
A student who was assigned to the baseline with no intents (V2) complained that the tutor would not let them use a solution entirely different from the expected one: “The AI only listens to key points that it wants, and ignores everything else (e.g. my idea of using […])” A good tutor should encourage the use of an original method, but the model considers it a mistake. This highlights a real pedagogical problem, as in this case, the tutor is harmful to the learning by restricting the student’s exploration. This insight shows that Pedagogical Steering is not just about telling or not the answer, but rather about a collection of different pedagogical properties LLMs lack.
6 Conclusion and Limitations
In this work, we identify the pedagogical limitations of LLMs in multi-turn conversational tutoring and present it as the Pedagogical Steering problem. We introduce the StratL algorithm to address this problem. StratL is based on Learning Sciences and utilizes a student state tracing and tutoring intent selection based on expert-defined transition graphs. We used the advanced but effective learning design Productive Failure in high school math to test it. A field study with high school students demonstrates faithfulness to the learning design and learner-perceived usefulness. We further release a dataset of 20 problems which could serve as a testbed for future improvements.
Evaluation aspects. The field study uses a rubric-based evaluation and learner-perceived usefulness in an end-to-end scenario. We argue additional evaluation metrics such as multiple human teachers evaluating the quality across the dimensions such as engaging the student or quality of questioning. Moreover, as we evaluated StratL through a field study, learners might have taken the exercise less seriously than they would for an exercise that is officially part of their curriculum. Further studies with increased ecological validity should be conducted.
Scalability. While tutoring can also be explored as a standalone experience, it is important to consider how systems like StratL could be integrated into a classroom or as part of a formal education context. As such, further research should explore (1) how to support teachers in adjusting and personalizing the tutor’s strategy, (2) how to provide relevant feedback to the teacher in order to ensure continuity with classroom content, and (3) how such tutors should be best integrated in existing pedagogical structures.
Social aspects. The social context of a learning experience plays an important role in learning, and “social surround” is even one of the three design layers of Productive Failure (Kapur and Bielaczyc 2012). Generally, supportive and coordinated relationships play an important role in learning (Roschelle and Teasley 1995), and research on how students perceive these tutors at the social level should be explored to paint a complete picture. As such, embodied agents could be considered (Sinha and Malhotra 2022). Beyond cognitive learning outcomes, it is also important to consider affective outcomes of learning activities, such as social-emotional competencies (Alzahrani, Alharbi, and Alodwani 2019), and how these are impacted by the interaction with digital tutors over human ones.
Data Privacy. The field study was approved by the university IRB (Proposal EK-2024-N-256). All data collected were anonymized and no sensitive student information was stored.
6.1 Acknowledgments
We are very grateful to the Math Department of the School of Science and Technology in Singapore for supporting and hosting our field study.
Appendix A Appendix
A.1 Taxonomy and Tutoring Strategy Modeling
We introduce a Learning-Sciences–based taxonomy of tutoring intents denoted $\mathcal{I}$ in Section 3.1. An intent is defined as the goal of a dialog utterance, such as correcting a mistake. The taxonomy groups all possible intents a tutor implemented with StratL can have during the tutoring session. We establish this taxonomy using the theories of Scaffolding (Reiser 2004), Affective Cognition (Lepper et al. 1993) and Productive Failure (Sinha and Kapur 2019; Kapur and Bielaczyc 2012).
We propose the following four categories of intents:
  1. 1.
    Scaffolding. The goal is to restrict the solution space, for example by decomposing the problem in steps, focusing the student’s attention on what is important, keeping track of what they did, or offloading unimportant parts of the task such as computations.
  2. 2.
    Problematizing. This category of intents aims at eliciting a deeper understanding through active engagement with the problem. In contrast to scaffolding, problematizing interventions expand the solution space by asking the student to generalize their answer, consider its limits, or understand its broader implications.
  3. 3.
    Affective. These intents aim at supporting the student in the often frustrating process of learning, and at drawing their interest.
  4. 4.
    Generic. The intents in this category serve broader conversational purposes, like greetings.
We subdivide and organize them in the taxonomy in Table 3.
Category Intent $I$ Example
Structuring (Reiser 2004) Guide Self-correction Take a second look at your formula.
Correct Are you sure AB is the hypotenuse?
Seek Strategy How would you proceed from now?
Hint Try to use the Pythagorean theorem.
State The Pythagorean theorem states that ….
Offload The result of this computation is ….
Problematizing (Reiser 2004) Identify Limits Can you think of a case where it fails?
Prompt Intuition/Hypothesis In your own words, what could be ….
Elicit Articulation Can you formally write this idea?
(Self-)Reflect (Katz, O’Donnell, and Kay 2000) What is the main takeaway of today?
Affective (Lepper et al. 1993; Sinha and Kapur 2019) Maintain Sense of Challenge I dare you to compute ….
Bolster Self-Confidence Seems like you outsmarted me!
Promote Sense of Control How do you want to call this triangle?
Evoke Curiosity In engineering, these equations are used ….
Generic (Macina et al. 2023a) State Greetings/Farewell Hello! How can I help you today?
Other Yes, you are allowed to use a calculator.
Table 3: Learning Sciences–based taxonomy of tutor intents.
A.2 Implementation
Student State Tracing
The set $\mathcal{S}$ (as per the notation used in Section 3.2) of labels (also called features) we choose from during the Student State Tracing procedure have been discussed with learning scientists and the prompt structure follows one used by the CLASS framework (Sonkar et al. 2023). We implement the assessment as a multi-label natural language classification task: the last student’s utterance is mapped to a predefined set of features representing the student’s intents, errors’ type, and emotional state. The full list of features can be read from the prompt in Figure 3 (letters a–m in the prompt). We call GPT-4o version ‘gpt-4o-2024-08-06’ on this prompt, with a temperature of 0 and a top_p parameter of $0.1$.
1A student and their tutor are working on a math problem:
2*Problem Statement*:
3##
4{pb}
5##
6
7The *Provided Solution* of this problem is:
8##
9{sol}
10##
11
12The tutors utterances are preceded by "Tutor:" and the students utterances are preceded by "Student:".
13
14Analyze the last students utterance.
15select all the feedbacks that apply from "a,b,c,d,e,f,g,h,i,j,k":
16
17a) The student is using or suggesting a wrong method or taking a wrong path to solve the problem
18b) The student made an error in the algebraic manipulation
19c) The student made a numerical error
20d) The student provided an intuitive or incomplete solution
21e) The students answer is not clear or ambiguous
22f) The student correctly answered the tutors previous question
23g) The student is explicitly asking about how to solve the problem
24h) The student is explicitly asking the tutor to state a specific theorem
25i) The student is explicitly asking the tutor to do a numerical calculation
26j) The student and tutor arrived at a complete solution for the entirety of the initial *Problem Statement*
27k) The student and tutor arrived at a complete solution for the entirety of the initial *Problem Statement* equivalent to the method provided in the *Provided Solution*
28
29Moreover select if relevant some emotional states from "l,m":
30l) The student shows a strong lack of motivation
31m) The student shows a strong lack of self-confidence
32
33Proceed step by step. First briefly justify your selection, then provide a string containing the selected letters.
34Answer in the following json format:
35##
36{{
37 "justification": "..",
38 "selection": ".."
39
40}}
41##
42Analyze the last students utterance.
43{
Figure 3: Student State Tracing Prompt to select the student state features, from those enumerated with lowercase letters (a, …, m).
We require the model to generate a justification for its assessment before writing the final selection. This justification serves a double purpose. First, it enhances the system’s explainability. We used these justifications to understand why the LLM selected some erroneous labels and used them to incrementally improve the system, including the formulation of the prompt. Second, this technique usually enhances the accuracy of classification task (Wu, Terry, and Cai 2021). For this purpose, it is important to require the justification to be written before the final label selection, as we use an autoregressive LLM, which can only consider context monodirectionally from the left. In the prompt, we use ‘##’ as a delimiter between the instructions and grounding information (such as the problem and its solution), as suggested in (OpenAi 2024). In lines 2 and 7 we surround “Problem Statement” and “Provided Solution” by ‘*’ to facilitate their reference later in the prompt (Reynolds and McDonell 2021). The core selection prompt from lines 15 to 31 is inspired by the assessment used in the CLASS framework (Sonkar et al. 2023). The exact formulation for each label has been optimized by trial and error over development iterations of the prototype. Lines 33 and 42 summarize the goal and ask the LLM to work step-by-step, a prompting technique drastically improving performance on reasoning benchmarks (Kojima et al. 2022). The last line, containing the first character of the expected answer (‘{’), acts as a syntactical constraint to guide the LLM to follow the JSON format (Reynolds and McDonell 2021).
The LLM also keeps in history the past utterances and their corresponding assessments.
A.3 Intent Selection
As detailed in Section 3.3 The Intent Selection is the procedure by which the tutor transitions from its previous intents $\mathbf{i}_{n}$ to its next $\mathbf{i}_{n+1}$ conditioned on the student state $\mathbf{s}_{n}$. This transition is represented by a deterministic transition graph. We recall that multiple intents can be selected. To transition from $\mathbf{i}_{n}$ to $\mathbf{i}_{n+1}$. One can follow the following algorithm:
  1. 1.
    For each intent $I$ in $\mathbf{i}_{n}$, start in the corresponding node in the transition graph.
  2. 2.
    The transitions from one node to the other(s) are conditioned on $\mathbf{s}_{n}$. If a transition condition (written on the arrows in the graph) is satisfied by $\mathbf{s}_{n}$, then follow the arrow and add to $\mathbf{i}_{n+1}$ the intent corresponding to the new node reached. Multiple arrows can be followed at the same time, and every corresponding intent is added. If no condition is met, no intent is added ($I$ is not either).
  3. 3.
    Repeat step 2 for each intent $I$ in $\mathbf{i}_{n}$.
For our implementation, we use an Intent Selection defined in collaboration with Productive Failure domain experts. We represent our algorithm by the transition graph in Figure 4. Each node of the graph represents a state, corresponding to an intent of Table 3. Arrows correspond to transitions between states. The letters used in the logical conditions on the arrows correspond to the Student State Tracing features detailed in Figure 3. For example, ‘b’ corresponds to the state feature “The student made an error in the algebraic manipulation”. These labels are combined in logical statements that condition the transitions. Arrows with no starting node represent transitions that can be taken from any state. We represent them this way for the sake of clarity. We recall that we allow for multiple intents, so if multiple transitions are possible at step $t$, they are all followed, and all corresponding intents are considered at step $t+1$. The Learning Science explanations for this graph are detailed in Appendix A.5.
Category Intent Text of the prompt addition
Structuring Guide Self-correction Make the student identify by themself the error in their answer.
Correct Correct the student’s mistake by giving them a hint.
Seek Strategy Encourage and make the student find on their own what is the next step to solve the problem, for example by asking a question.
Hint Give a hint to the student to help them find the next step. Do *not* provide the answer.
State State the theorem or definition the student is asking about.
Offload Correct and perform the numerical computation for the student.
Problematizing Identify Limits Make the student identify the limits of their reasoning or answer by asking them questions.
Prompt Intuition/Hypothesis Ask the student start by providing a guess or explain their intuition of the problem.
Elicit Articulation Ask the student to write their intuition mathematically or detail their answer.
(Self-)Reflect Step back and reflect on the solution. Ask to recapitulate and *briefly* underline more general implications and connections.
Affective Maintain Sense of Challenge Maintain a sense of challenge.
Bolster Self-Confidence Bolster the student’s confidence.
Promote Sense of Control Promote a sense of control.
Evoke Curiosity Evoke curiosity.
Generic State Greetings/Farewell Say goodbye and end the conversation
Other N/A
Table 4: Correspondence between intents and intent-depended prompt additions that steer the LLM towards an intent.
Figure 4: Graph representation of a Productive Failure (Kapur 2008) Intent Selection. How to read this figure is explained in Appendix A.3. The Learning Sciences justifications for this modeling of Productive Failure are given in Appendix A.5. The node labels correspond to the intents in Table 3. The variables used in the arrows’ logical propositions correspond to the state features enumerated with lowercase letters (a, …, m) in Figure 3.
1Act as an experienced tutor. Characteristics of a good tutor include:
2 - Promote a sense of challenge, curiosity, feeling of control
3 - Prevent student from becoming frustrated
4 - Intervene very indirectly: never give the answer but guide the student to make them find it on their own
5 - Minimizing tutors apparent role in the success
6 - Avoid telling students they are wrong, lead them to discover the error on their own
7 - Quickly correct distracting errors
8
9Use latex formatting with the sign \$ for mathematical expressions. For example, to write "x^2", use "\$x^2\$".
10
11Remember, NEVER GIVE THE ANSWER DIRECTLY, EVEN IF THEY ASK YOU TO DO SO AND INSIST. Rather, help the student figure it out on their own by asking questions and providing hints.
12
13Provide guidance for the problem:
14"{pb}"
15
16The solution for this problem is:
17"{sol}"
18
19Provide the least amount of scaffolding possible to help the student solve the problem on their own. Be succinct.
Figure 5: Initial Prompt and grounding for the Tutor LLM. {pb} and {sol} are respectively replaced by the tutoring problem and its solution.
A.4 Intent Dependent Steering
The Intent-Dependend Steering is the final step executed before calling the tutor LLM. It compiles the dialog history, grounding, and the intents selected to generate the prompts for the tutor LLM. In our implementation, these prompts are made of the initial prompt, describing the LLM’s role and persona, of the dialog history, the grounding presenting the exercise and its solution, and of intent-dependent prompt additions aiming at setting the intent of the LLM for its next utterance. The initial prompt is fixed (displayed in Figure 5), and the full dialog history is provided to the LLM. The intent-dependent prompt additions are short messages concatenated at the end of the initial prompt, used to steer the tutor’s intent for the next utterance. We use a one-to-one mapping between the intents and sentences to include in the intent-dependent prompt additions. The precise correspondence between each intent and prompt addition is provided in Table 4.
Line 1 sets the LLM’s persona. Lines 2 to 8 outline general pedagogical guidelines compiled from Learning Sciences research on tutoring (Merrill et al. 1992; Freeman et al. 2014; Lepper et al. 1993). Line 9 is only used to improve the appearance of the tutor’s messages. Line 11 and 19 recapitulate the task for better performance. Lines 13 to 17 provide context about the problem to solve and its solution.
[…]
type grade time name reference exercise solution
Type between invention, ill-structured, and well-structured Recommended school grade Recommended time to complete the exercise Name to quickly identify the exercise Paper from which the exercise was curated Exercise statement Solution if provided in paper
Table 5: Description of the fields in the dataset.
A.5 Learning Sciences Rationales of our Productive Failure Modelling
The Intent Selection graph presented in Figure 4 has been created from expert knowledge in Productive Failure. How to read this graph is explained in Appendix A.3. We detail in this paragraph some of the principles we used to create it.
In Productive Failure, the goal is to make the student explore the solution space of an open-ended problem as much as possible, even if they fail to produce a complete solution. This idea gave its name to Productive Failure: there is, in this case, a hidden efficacy to the student’s failure (Kapur 2008). To this aim, in Figure 4 any mistake (features a, b or c in Figure 3) leads to a Guide Self Correction tutoring intent. The goal is to make the student spot and correct their mistake by themself. If the error persists, we use the intent Correct that gives a hint to the student about their mistake, but no intent seeks to directly correct it. If the mistake is computational only, we use the intent Offload that prompts the LLM to perform the computation. This allows the student to focus on exploring the solution space rather than wasting time on computations. For the same reason, if the student asks for a basic theorem or definition they forgot (feature h), we use the intent State to answer their question and help them focus on the reasoning.
A good tutor should intervene very indirectly (Merrill et al. 1992): a good principle to follow as a tutor is to give as little scaffolding as one can get away with. The intent used whenever a student manages to correct their mistake (arrow labeled f leaving the Guide Self-Correction node in Figure 4) or when the student suggests an intuitive or incomplete solution (state feature d) is the Seek Strategy intent. This encourages the student to pursue their line of thought without heavier scaffolding.
One of the challenges of Productive Failure is to create a classroom culture that allows for failure and promotes psychological safety. Indeed, as students deal with open-ended problems before being given the tools to solve them, the experience can be stressful. In traditional classroom settings, the focus is directed towards getting the right solution and students are often evaluated on this with a grade. One of the most crucial design features of a Productive Failure problem session is to shift this goal from solving the problem to generating multiple RSMs regardless of their correctness (Kapur and Bielaczyc 2012). As explained in Kapur and Bielaczyc (2012), “students are used to seeking assistance from their teachers so much so that they do so even before sufficiently trying to solve problems themselves. At the same time, teachers are just as used to providing assistance when it is asked for so much so that often opportunities for students to generate and explore RSMs are missed”. For this reason, in our modelling, if the student explicitly asks for help the intent Prompt Intuition/Hypothesis is triggered. This intent encourages the student to start with a guess or an intuition. This pushes students to try with no fear of being incorrect. Once the student is successful in providing an intuitive solution, or if they confidently propose an incomplete one (feature e), then the Elicit Articulation intent is used to encourage an RSM creation by formulating their idea in concrete math. To mitigate stress, that can hinder the student’s learning (Lepper et al. 1993), we trigger a Hint intent to ease the student’s progress if they show signs of lack of motivation or self-confidence.
A good Productive Failure problem allows for multiple solutions (and often a preferred one, called canonical solution). Then, if the student manages to propose a non-canonical solution to the problem (feature j), they should be encouraged to continue exploring. The Identify Limits intent is triggered to ask the student to identify the limitations of their solution, encouraging them to develop a new one.
If the student reaches the canonical solution (feature k), we use the intent (Self-)Reflect. It has been proven that it is beneficial from a learning point of view to engage in self-reflection (i.e. looking back critically at the work done) after a problem session (Katz, O’Donnell, and Kay 2000). Following this turn, the tutor is made ready to end the conversation through the State Greetings/Farewell intent.
A.6 Dataset
We collected 20 problems from the Productive Failure literature  (Kapur 2008; Kapur and Bielaczyc 2012; Kapur 2010; Kerrigan, Weber, and Chinn 2021; Kerrigan 2018). We present the columns of the dataset in Table 5. We curated three types of problems: Invention problems ask the student to invent a new way to solve a problem. They typically admit several correct but suboptimal answers, as long as the student can justify them. Students are not expected to have previous exposure to the tools canonically used to solve the exercise. The problem Consistency is a good example of invention problem. Ill-structured problems are open-ended problems that contain some non-relevant information and a convoluted story. These problems often admit a single solution, but there can be several ways to reach it. The problem country is an example of such a problem. Well-structured problems are more direct versions of the ill-structured problems. They resemble classical exercises that students usually solve in a mathematics class. More information about the different types of exercise in Productive Failure and their role can be found in (Kapur 2008) and (Kapur and Bielaczyc 2012).
A.7 Test Problems
Figures 6 and 7 display the two test problems we used in our field test. They are also included in the dataset we release.
A.8 PF Score Details
We present in the following paragraphs the precise reasoning sub-steps used for the Productive Failure Score grading as explained in Section 4.4.
For problem Consistency, we define these reasoning steps as 1. Giving an intuitive definition of consistency related to the variation of scores, 2. Thinking about comparing the scores to a baseline (central value or previous year), 3. Considering absolute differences (by taking the absolute value or squaring) since only the magnitude of the change matters, and 4. Thinking about a way of aggregating all the differences into a single score, for example by summing or averaging them.
For problem Country, we define these steps as 1. Thinking about using a ‘standard divisor’ which represents the number of seats per inhabitant, 2. Thinking about rounding down the number of seats not to attribute more seats than the total, and 3. Thinking about using the remainders to distribute the additional seats. We linearly scale the PF score to 4 in this case.
A new country has recently been founded. The country is split into six states, call them A, B, C, D, E, and F.
The population of state A is 1,646,000 people, the population of state B is 6,936,000 people, the population of state C is 154,000 people, the population of state D is 2,091,000 people, the population of state E is 685,000 people, and the population of state F is 988,000 people.
There are 250 seats available on a legislative body to govern the new country. How many seats should be assigned to each state so that each state would receive a fair representation?
Show your work and justify why you think your method is correct.
Solution:
We assign sits proportionally to the population of each state. Since the results of the divisions are not integers, we round down the number and then distribute the remaining sits to stats having the largest remainders.
Total population = 1,646,000 (A) + 6,936,000 (B) + 154,000 (C) + 2,091,000 (D) + 685,000 (E) + 988,000 (F) = 12,500,000
Standard divisor = Total population / Number of seats = 12,500,000 / 250 = 50,000
Initial quotas:
- A: 1,646,000 / 50,000 = 32.92 → 32 seats
- B: 6,936,000 / 50,000 = 138.72 → 138 seats
- C: 154,000 / 50,000 = 3.08 → 3 seats
- D: 2,091,000 / 50,000 = 41.82 → 41 seats
- E: 685,000 / 50,000 = 13.70 → 13 seats
- F: 988,000 / 50,000 = 19.76 → 19 seats
Total initial seats assigned = 32 + 138 + 3 + 41 + 13 + 19 = 246
Seats left to distribute = 250 - 246 = 4
Distribute the surplus seats based on largest remainders:
Remainders (from the divisions above):
- A: 0.92
- B: 0.72
- C: 0.08
- D: 0.82
- E: 0.70
- F: 0.76
The four highest remainders are from states A, B, D, and F. Give one extra seat to each.
Figure 6: Problem Country and its solution.
The organizers of the Premier League Federation have to decide which one of the three players Mike Arwen, Dave Backhand and Ivan Right - should receive the ”The Most Consistent Player for the Past 5 Years” award. Table 1 shows the number of goals that each striker scored between 2019 and 2023.
The organizers agreed to approach this decision mathematically by designing a measure of consistency. They decided to get your help. Here is what you must do:
(1) Design as many different measures of consistency as you can.
(2) Your measure of consistency should make use of all data points in the table.
Table 1. Number of goals scored by the three players in the Premier League between 2019 and 2023.
Year Mike Arwen Dave Backhand Ivan Right
2007 13 12 14
2008 12 14 10
2009 15 16 18
2010 17 15 18
2011 13 13 15
Solution:
The concept of variance and standard deviation is unknown to students. Any measure proposed by the student is acceptable as long as it can be justified to measure consistency. The goal is for them to construct their own measure of consistency and justify it based on the data provided.
Example of canonical solution: computing the variance (or standard deviation) for each player (standard deviation is also valid):
First, compute the mean:
Mean number of goals for Mike: 14
Mean number of goals for Dave: 14
Mean number of goals for Ivan: 15
Then, compute the sum of square deviations from the mean for each player:
Sum squared deviation for Mike: 16
Sum squared deviation for Dave: 10
Sum squared deviation for Ivan: 44
Then divide by the number of data points to get the variance:
Variance for Mike: 12/5 = 3.2
Variance for Dave: 10/5 = 2
Variance for Ivan: 44/5 = 8.8
So according to the variance, Dave is the most consistent player.
Figure 7: Problem Consistency and its solution.
A.9 Student State Tracing Evaluation
As explained in Section 5.1, we evaluated the accuracy of our Student State Tracing procedure, including a comparison against baselines.
Table 6: Test micro-averages for different versions of the Student State Tracing.
$\mu$-average precision $\mu$-average recall $\mu$-average F1
Full version 0.82 0.75 0.77
No Justification 0.71 0.72 0.70
Short Memory 0.64 0.75 0.68
Methodology
We test the procedure using two tutoring conversations of about ten dialog turns for each of the two test problems detailed in Appendix A.7. These conversations are manually annotated to obtain the reference data. As detailed in Appendix A.2, we use an LLM to perform this task in our implementation of StratL. This LLM is prompted with the full conversation history and is asked to output a justification for the labels it selected. We compare this implementation to a version not outputting justifications and a version that only keeps the three last utterances in the prompts’ history to increase the processing speed.
As a multi-label classification task, we compute for each label $i$ the precision, recall, and F1 score. We aggregate these metrics across labels using an average weighted by the relative cardinality of each label, known as the micro average ($\mu$-average). We report the results in Table 6.
Results
Our simple Student State Tracing already achieves satisfactory results. The use of justification indeed seems to increase the State Tracing’s performance as we can observe a higher precision and F1 score for the version using it compared to the version not justifying its answers. It is worth remembering that this increased performance comes at a higher cost, since the LLM has to generate additional tokens for the justification. The State Tracing model keeping in memory only the last three dialog pairs does not perform as well as the full version: it thus appears important to provide the whole context of the conversation to obtain reliable results.
A.10 V4 LLM Intent Selection Prompt
Figure 8 displays the prompt of the LLM used for the Intent Selection in the baseline V4.
1A student and a tutor are working on a math problem. The tutor responds to the students messages following some intents. Before responding, the tutor first selects the most appropriate intents from a *Taxonomy*.
2The tutor selects the intents based on the Productive Failure teaching method. The goal is to let the student explore the space of solution by themself to generate as many representation and solution methods as possible, even if they are only partially correct.
3Select the best set of intents for the tutors next message based on the following *Taxonomy*:
4
5*Taxonomy*:
6##
7 P_LIMITS : Ask the student to identify some limits of their reasoning or answer.
8 P_HYPOTHESIS : Ask the student to start by providing a guess, hypothesis or explain their intuition of the problem.
9 P_ARTICULATION : Ask the student how they can write their intuition mathematically or detail their answer.
10 P_REFLECTION : Reflect on the solution, recapitulate, and underline more general implications.
11 S_SELFCORRECTION : Ask a question to hint the student to identify errors in their answer.
12 S_CORRECTION : Correct a students mistake
13 S_STRATEGY : Make the student find and perform the next step to solve the problem without giving a hint.
14 S_State : State a theorem or definition asked by the student.
15 S_OFFLOAD : Perform a computation for the student
16 S_HINT : Give a hint to the student
17 G_GREETINGS : Say goodbye to the student
18##
19
20The tutors previous intents for their last message to the student were: {previous_intent}.
21The student answered the last tutors message. An assessment of their answer has been conducted using the following codes:
22
23assessment codes:
24##
25a) The student is using or suggesting a wrong method or taking a wrong path to solve the problem
26b) The student made an error in the algebraic manipulation
27c) The student made a numerical error
28d) The student provided an intuitive or incomplete solution
29e) The students answer is not clear or ambiguous
30f) The student correctly answered the tutors previous question
31g) The student is explicitly asking about how to solve the problem
32h) The student is explicitly asking the tutor to state a specific theorem
33i) The student is explicitly asking the tutor to do a numerical calculation
34j) The student and tutor arrived at a complete solution for the entirety of the initial *Problem Statement*
35k) The student and tutor arrived at a complete solution for the entirety of the initial *Problem Statement* equivalent to the method provided in the *Provided Solution*
36l) The student shows a strong lack of motivation
37m) The student shows a strong lack of self-confidence
38##
39
40The assessment for the last students message is: {assessment_codes}.
41
42Based on the assessment and the last tutors intents, select the most appropriate set of intents for the next tutors message.
43Proceed step by step. First, briefly justify your selection, then provide a list containing the selected intents from the *Taxonomy*.
44
45Answer in the following json format:
46##
47{
48 "justification": "...",
49 "intents": []
50
51}
52##
53{
Figure 8: Prompt for the LLM Intent Selection in V4.
A.11 User interface
Figure 9: User interface of our math tutor prototype used by students in the field study.